Find the evolute of

x

2

a2

y

2

b

2

= 1 as the envelope of the normals.


1
Expert's answer
2021-12-23T17:30:47-0500

"x^\n\n2\/\n\na^2\n\n+\n\ny^\n\n2\/\n\nb^\n\n2\n\n= 1"

"x=acost,y=bsint"


evolute of a curve:

"X(t)=x(t)-\frac{y'(t)(x'(t)^2+y'(t)^2)}{x'(t)y''(t)-x''(t)y'(t)}"


"Y(t)=y(t)+\frac{x'(t)(x'(t)^2+y'(t)^2)}{x'(t)y''(t)-x''(t)y'(t)}"


"X(t)=acost-\frac{bcost(b^2cos^2t+a^2sin^2t)}{absin^2t+abcos^2t}=acost-\frac{cost(b^2cos^2t+a^2sin^2t)}{a}"


"Y(t)=bsint-\frac{asint(b^2cos^2t+a^2sin^2t)}{absin^2t+abcos^2t}=bsint-\frac{sint(b^2cos^2t+a^2sin^2t)}{b}"


equation of normals ar point (x1,y1):

"(y-y_1)=-y'(x_1)(x-x_1)"

"a^2x\/x_1 - b^2y\/y_1 = a^2 - b^2"


then:

"a^2x\/(acost) - b^2y\/(bsint) = a^2 - b^2"

differentiate respect to t:

"xcost+ysint=(a^2-b^2)cos2t"